OpenStudy (anonymous):

Find all values of x for which the function is differentiable. y=cuberoot(3x-6)+5 Am I supposed to set the derivative equal to 0?? I have the derivative.. y'=(1/3)(3x-6)^(-2/3).. now what do I do with this?

myininaya (myininaya):

find the domain of y'

OpenStudy (anonymous):

The domain is all real numbers, right?

myininaya (myininaya):

\[y=\sqrt[3]{3x-6}+5=> y'=\frac{1}{3}(3x-6)^{\frac{1}{3}-\frac{3}{3}}+0\] \[y'=\frac{1}{3}(3x-6)^\frac{-2}{3}=\frac{1}{3(3x-6)^\frac{2}{3}}\] 3x-6 can't be zero, right?

OpenStudy (anonymous):

Ohh yes, I forgot about the negative exponent. So I find when 3(3x-6)^(2/3) = 0?

myininaya (myininaya):

find the domain of y'

myininaya (myininaya):

all real numbers except when 3x-6=0 which is when x=2

myininaya (myininaya):

all real numbers except x=2